OpenStudy (anonymous):

Suppose the conducting wire (with resistance R) was replace with one of the same material but has its dimension doubled (the length and the radius having twice their original value. What is the new resistance in comparison to the previous resistance?

hartnn (hartnn):

if you use the formula \(\large R=\rho \frac{L}{A}=\rho\frac{L}{\pi r^2}\) its given that \(\rho \) remains same \(L'=2L,r'=2r\) so \(\large new \: R=R'=\rho \frac{2L}{\pi (2r)^2}=....?\) so, can u find R' in terms of R ?